# GraphCompNet: A Position-Aware Model for Predicting and Compensating Shape Deviations in 3D Printing

### v1

@@ Abstract
Shape deviation modeling and compensation in additive manufacturing (AM) are pivotal for achieving high geometric accuracy and enabling industrial-scale production. While traditional analytical and statistical methods laid the foundation, recent advancements in machine learning (ML) have improved prediction and compensation precision. However, critical challenges persist, including generalizability across complex geometries and adaptability to position-dependent variations in batch production. Traditional methods of controlling geometric deviations often rely on complex parameterized models and repetitive metrology, which can be time-consuming yet not applicable for batch production. In this paper, we present a novel approach to address this challenge of ensuring geometric precision and accuracy in position-dependent powder bed fusion (PBF) production. The proposed GraphCompNet presents a novel computational framework integrating graph-based neural networks with a generative adversarial network (GAN)-inspired training paradigm. The framework leverages point cloud representations and dynamic graph convolutional neural networks (DGCNNs) to model intricate geometries while incorporating position-specific thermal and mechanical variations. A two-stage adversarial training process iteratively refines compensated designs using a compensator-predictor architecture, enabling real-time feedback and optimization. Experimental validation across various shapes and positions demonstrates the framework’s ability to predict deviations in freeform geometries and adapt to position-dependent batch production conditions, significantly improving compensation accuracy (35% to 65%) across the entire printing space, addressing position-dependent variabilities within the print chamber. The proposed method advances the development of a Digital Twin for AM, offering scalable, real-time monitoring and compensation capabilities. This work bridges critical gaps in AM process control, paving the way for high-precision, automated, and industrial-scale design and manufacturing systems.
Index Terms:
Additive Manufacturing, Powder-bed Fusion, Shape Deviation, Shape Compensation, Graph Neural Network, Quality control, Digital Twin
publicationid: pubid: $0$–$00$ 2021 IEEE
@@ I Introduction
Powder Bed Fusion (PBF) is an advanced additive manufacturing (AM) technique that fuses powdered materials—typically metals, plastics, or ceramics—using a laser or electron beam. By precisely controlling energy deposition at the voxel level, PBF creates highly complex and detailed structures with fine resolution (80-250 microns). It is widely used in industries like aerospace and biomedicine for high-quality output and the ability to produce multiple parts in a single build. Among PBF methods, Multi Jet Fusion (MJF) [1] improves production efficiency by injecting a rapid agent at the pixel level and projecting thermal energy across the build area, enabling faster print speeds without compromising accuracy or quality. A critical challenge in AM in general, however, is achieving consistent geometric accuracy.
@@
For instance, as shown in Fig. 1, a molded fiber part placed at the center of the print chamber (blue in Fig. 1) demonstrates a maximum deviation of 6 mm from the nominal design, with a standard deviation of 1 mm. In contrast, a part placed at the bottom of the chamber (pink in Fig. 1) shows a maximum deviation of 7 mm and a standard deviation of 1.1 mm. These parts display distinct deviation patterns across their geometries, as highlighted by the varying color map distributions. Achieving high precision in AM is challenging due to the complex interactions of multiple factors, including material properties, part orientation, printer settings, calibration, layer-by-layer deposition, energy distribution, temperature gradients, and residual stresses, among others. Early research focused on analytical and statistical models to predict and mitigate these deviations, but their reliance on simplified assumptions limited applicability to real-world geometries. Recent advances in computational methods, particularly machine learning (ML), have introduced data-driven models that enhance prediction accuracy and scalability. Despite progress, critical challenges hinder the transition of these models to industrial-scale AM.
Existing deviation prediction models face challenges with generalizing across complex geometries. FEA-based models [2] are computationally expensive for real-time use, while data-driven frameworks for 3D deviation prediction are often validated only for simple shapes (e.g., domes). Furthermore, compensation models frequently overlook position-dependent variations inherent in batch production (referred to as ”position-aware” below), such as the significant deviations caused by spatial factors like thermal gradients. Additionally, the lack of experimental validation in prior research limits the practical applicability of these models in industrial settings.
Since batch production is essential for industrial applications, overcoming these challenges is crucial for developing more efficient, automated systems like a Digital Twin, which can improve precision, control, and scalability in AM. A breakthrough in this area would significantly enhance both design accuracy and mass production reliability, driving further innovation in the field. The remainder of the paper is structured as follows: Section II reviews deviation prediction and compensation models, as well as the research challenges, Section III details the proposed methodology, Section IV presents experimental results, and Section V concludes with future directions.
@@ II State-of-the-art and research challenges
Shape deviation modeling and compensation in AM have been critical for improving geometric accuracy and quality control. Initially, research focused on analytical and statistical models, recent advancements have incorporated computational techniques such as machine learning approaches. This section highlights three key areas: (1) review of Deviation Prediction Modeling, (2) review of Deviation Compensation Modeling, and (3) Research Challenges and Our Proposed Methodology, emphasizing the transition to data-driven strategies and the hurdles in achieving industrial-scale precision in AM mass production.
@@ II-A Deviation Prediction Modeling
Early research on shape deviation often employed statistical and analytical models for parameterization, both in 2Dand 3D [3, 4, 5]. Schmutzler et al. [6] innovated by framing shape deviation prediction as a non-rigid shape registration problem. This approach, utilizing B-spline-based registration, predicted deviations and compensated for CAD parts by shifting CAD vertices in the opposite direction of the learned deviation [7]. Huang et al. [8] developed statistical models using Polar and Spherical Coordinate Systems (PCS, SCS). These models address predicting in-plane and out-of-plane deviations, though their reliance on experimental setups and calibration limited scalability and adaptability. Additionally, finite element analysis (FEA) was applied to simulate part deformation based on real process instructions .
More recently, machine learning (ML) techniques have been incorporated into deviation prediction to improve accuracy. Ferreira et al. [9] used Bayesian neural networks to model and compensate for deviations. Decker et al. [10] developed a random forest model for 3D freeform shapes, enabling rapid, minimal human intervention. Convolutional neural networks (CNNs) [11] and Shape Deviation Generators [12] have been applied to detect complex deviation patterns, while Wang et al. [13] extended ML to both smooth and non-smooth shapes. Li et al. [14] proposed in-situ deviation monitoring during printing.
Despite advances, several challenges remain. Analytical and statistical models are computationally expensive, with analyses on a single part often taking hours. While ML methods show promise, they are often limited to simpler geometries (such as simple freeform shapes based on cylinders and polyhedrons), leaving a gap for more intricate 3D geometries. This highlights the need for more efficient and generalized models capable of addressing both computational constraints and the complexity of real-world AM geometries.
@@ II-B Deviation Compensation Modeling
[…]
To address these challenges, advanced algorithms have been introduced. Bayesian models [9] use uncertainty quantification to correct for geometric errors in stereolithography printing, improving robustness across different printer setups. Decker et al. [10] applied random forest models to predict deviations in fused deposition modeling (FDM) with a small training dataset, demonstrating over 44% shape deviation. Though domain knowledge based on the geometry is essential to define the predictor variables for the random forest model. Hong et al. [19] used artificial neural networks (ANN) to enhance dimensional accuracy in truss lattice structures, outperforming traditional methods in selective laser melting. The method highlighted the potential of using a data-driven approach to generate flexible, free-form geometries, providing enhanced compensation. Despite recent advancements, existing compensation techniques generally rely on shape deviation models for isolated individual shapes. The substantial effect of thermal processes on shape deviations requires modeling parts in different positions within the printer, which demands considerable resources for creating position-specific deviation models, even for a single geometry.
@@ II-C Research Challenges and Proposed Methodology
Despite significant advancements in shape deviation modeling and compensation, key challenges persist, particularly regarding scalability, data dependency, and generalization. These challenges emphasize the need for novel approaches that address these gaps, ultimately enhancing the efficiency and accuracy of AM. This research proposes a computational framework to predict and compensate for print geometry, incorporating position-specific variables such as energy differences in the print chamber of the same print geometry. Limited research has focused on addressing position-specific compensation in AM, making this contribution vital. The primary challenges in this context are:
@@
  • Generalizability Across Complex Printing Geometries: Exisiting works demonstrate modeling, predicting and compensating for 2D shape deviations [20, 21]. However, modeling and control 3D geometries taking into account the layer-wise interactions presents significantly greater challenges. Huang et al. [12, 13] introduced a data-analytical framework for 3D deviation, but its validation has been limited to simple shapes such as domes and stacked rectangles. Approximating non-smooth 3D freeform shapes using a combination of patches remains constrained by resolution limits and requires additional steps to subtract polyhedral shapes. Additionally, Finite Element Analysis (FEA)-based models are computationally expensive and impractical for complex industrial geometries and batch processes [22]. More recent methods using Convolutional Neural Networks (CNNs) [23, 11] for error prediction. Shen et al. [24] employed an Encoder-Decoder architecture to predict and compensate binary image data. However, those methods are limited by their resolution and lack of detail, due to preprocessing in fixed-size images or binary format.
  • Adaptability to Position-Dependent Batch Production: Traditional and machine learning-based compensation models often fail to consider the dynamic thermal and mechanical factors that influence AM, leading to suboptimal results under varying conditions [25]. Thermal-induced deformations significantly impact part distortion. Research by Chen et al. [26] shows how thermal variations across the print bed result in significant differences in part quality. For example, Figure 2 illustrates a dataset consisting of 140 identical bars placed at different positions within the printing bed. The shape deviation patterns exhibit significant variation, with parts positioned at the edges of the print chamber, such as Part ID 140 and Part ID 31 in Fig. 2 (b), experiencing greater deviations compared to those located at the center, such as Part ID 72 and Part ID 24. Hartmann et al. [27] highlighted the importance of batch production in industrial settings, demonstrating a data-driven compensation approach applied to stacked fin-shaped parts. However, their method is constrained by the need for iterative redesigns and was validated using only a single geometry.
@@
      DGCNN (Dynamic Graph Convolutional Neural Network): To further enhance the model’s ability to handle the complexities of point cloud data, we incorporate DGCNN as the backbone architecture. DGCNN dynamically constructs graph structures by connecting points based on local geometric relationships, allowing it to effectively capture the varying densities and topologies inherent in complex shapes. The EdgeConv operation within DGCNN adapts to these local variations, facilitating the aggregation of point features and improving the model’s capacity to identify fine-grained spatial patterns. This enables the model to achieve higher performance in tasks that require precise understanding of local geometry and complex spatial relationships, crucial for accurate shape prediction and deviation analysis in diverse 3D structures.
@@
      Incorporation of Part Position Information: First, we incorporate part position information into the input geometries. For parts translated along the x and y axes, we adjust the corresponding point cloud coordinates to account for the scale. For parts that are rotated, we apply a rotation matrix to the point cloud coordinates. As a result, the input data integrates both the part geometry and placement information. This enables the model to learn position-specific variables during training, allowing it to adapt to variations in print conditions and ensuring accurate predictions, even in the presence of environmental or process-related differences.
@@ III Methods
Review of Geometric Deep Learning. Convolutional neural networks (CNNs) have spearheaded significant advancements in computer vision and natural language processing However, the inherent non-Euclidean nature of 3D designs presents a challenge for directly implementing standard CNNs on CAD parts. Graph neural networks (GNNs) have garnered increasing attention for geometric applications, offering a framework that accommodates non-Euclidean data structures. In GNNs, graph structures $G=\{\mathcal{V},\mathcal{E}\}$, where $\mathcal{V}$ and $\mathcal{E}$ represent the sets of vertices and edges, are leveraged to define convolution operations tailored to the given data.
Graph neural networks can be broadly categorized into two families: spectral and spatial. The spectral family of GNNs is based on the concept of generalized Fourier transforms, utilizing eigen-decomposition of the graph Laplacian [28] for convolution operations. However, computing the graph Laplacian can be computationally expensive. To address this challenge, alternative approaches have been proposed, including approximating spectral filters using polynomial filters [29]. Spatial approaches define convolution operations $h_{\theta}$ with regard to edge information between vertices $e_{ij}\in\mathcal{E}$, offering a computationally efficient alternative [30]:
@@
where $l$ denotes $l$-th convolutions, $\theta$ is a set of learnable parameters. As a representation, Edge Convolution (EdgeConv) concatenates and aggregates features from $x_{i}^{(l)}$ and $x_{i}^{l}-x_{j}^{(l)}$, representing edge information with relative distance between neighboring vertices [31]. MoNet [30] projects neighboring points onto a pseudo-coordinate plane, then define convolution operations on projected points on the plane. Overall, the evolution of graph neural networks presents a promising path for tackling geometric applications, offering flexibility and scalability in handling non-Euclidean data structures inherent in 3D designs [32].
@@
Here, $G$ and $D$ represent the generator and discriminator, respectively. While the optimization problem above is equivalent to minimizing the Jensen-Shannon divergence (JSD) [33] when the discriminator is perfectly trained, JSD often leads to unstable GAN training, especially in the presence of singular measures [34]. To address this issue, Wasserstein-GAN was introduced, replacing JSD with the Wasserstein metric, which compares transportation costs between distributions $\mathbb{P}_{G}$ and $\mathbb{P}_{r}$. The 1-Wasserstein distance derived from Kantorovich-Rubenstein duality is employed for GAN optimization, instead of directly computing transportation mappings. To ensure the computation of 1-Wasserstein distance, both the generator and discriminator must exhibit 1-Lipschitz continuity during training. This can be achieved through techniques such as weight clipping [34] or gradient penalty.

### v3

@@ Abstract
Shape deviation modeling and compensation in additive manufacturing (AM) are pivotal for achieving high geometric accuracy and enabling industrial-scale production. While traditional analytical and statistical methods laid the foundation, recent advancements in machine learning (ML) have improved prediction and compensation precision. However, critical challenges persist, including generalizability across complex geometries and adaptability to position-dependent variations in batch production. Traditional methods of controlling geometric deviations often rely on complex parameterized models and repetitive metrology, which can be time-consuming yet not applicable for batch production. In this paper, we present a novel, process-agnostic approach to address the challenge of ensuring geometric precision and accuracy in position-dependent AM production. The proposed GraphCompNet presents a novel computational framework integrating graph-based neural networks with a generative adversarial network (GAN)-inspired training paradigm. The framework leverages point cloud representations and dynamic graph convolutional neural networks (DGCNNs) to model intricate geometries while incorporating position-specific thermal and mechanical variations. A two-stage adversarial training process iteratively refines compensated designs using a compensator-predictor architecture, enabling real-time feedback and optimization. Experimental validation across various shapes and positions demonstrates the framework’s ability to predict deviations in freeform geometries and adapt to position-dependent batch production conditions, significantly improving compensation accuracy (35% to 65%) across the entire printing space, addressing position-dependent variabilities within the print chamber. The proposed method advances the development of a Digital Twin for AM, offering scalable, real-time monitoring and compensation capabilities. This work bridges critical gaps in AM process control, paving the way for high-precision, automated, and industrial-scale design and manufacturing systems.
publicationid: pubid: 0000–0000/00\$00.00 ©  IEEE 2025
@@ I Introduction
Additive manufacturing (AM), commonly known as 3D printing, encompasses a diverse set of layer-by-layer fabrication processes capable of producing complex geometries with high dimensional precision. Among these, Powder Bed Fusion (PBF)—including Laser Powder Bed Fusion (L-PBF) and Multi Jet Fusion (MJF) [mjf-bio]—has emerged as a leading platform due to its fine spatial resolution (typically 80–250 microns) and ability to batch-produce intricate, high-performance components. Despite these advantages, PBF and other AM processes face persistent challenges in geometric fidelity, such as thermal distortion, shape deviations, and position-dependent variability across the build platform. These issues are common across AM methods and remain critical bottlenecks for industrial-scale deployment.
For example, Fig. 1 illustrates geometric deviations for two vertically stacked parts printed simultaneously in a multi-part build: Part 1 (Center - Blue) and Part 2 (Bottom - Pink). The heatmaps reveal distinct distortion patterns for each part, highlighting position-dependent deformation within the same build. Part 1 exhibits complex warping with edge uplift and non-uniform shrinkage, whereas Part 2 shows a more symmetric dome-like distortion with fewer localized defects. Such differences likely stem from variations in thermal boundary conditions and support interactions during printing.
@@
Achieving high precision in AM is challenging due to the complex interactions of multiple factors, including material properties, part orientation, printer settings, calibration, layer-by-layer deposition, energy distribution, temperature gradients, and residual stresses, among others. Early research focused on analytical and statistical models to predict and mitigate these deviations, but their reliance on simplified assumptions limited applicability to real-world geometries. Recent advances in computational methods, particularly machine learning (ML), have introduced data-driven models that enhance prediction accuracy and scalability. Despite progress, critical challenges hinder the transition of these models to industrial-scale AM.
Current deviation prediction models often struggle to generalize across complex geometries. Finite element analysis (FEA) methods [10.1115/1.4041626] are computationally intensive and unsuitable for real-time applications, while many data-driven approaches have only been validated on simple shapes (e.g., domes). Moreover, most compensation models do not explicitly consider position-dependent variations (referred to here as ”position-aware”), such as thermal gradients across the build plate. The lack of extensive experimental validation further limits practical deployment.
Because batch production is essential for industrial AM, addressing these challenges is critical for enabling automated systems like Digital Twins that enhance precision, control, and scalability. Such advancements would substantially improve design fidelity and production reliability, accelerating innovation in digital manufacturing.
This paper proposes a generalizable framework for modeling and compensating geometric deviations in AM, using PBF as the primary use case. The approach avoids reliance on process-specific assumptions, making it applicable to other AM platforms experiencing similar spatial and thermal effects. The rest of the paper is organized as follows: Section II reviews related work in geometric deviation prediction and compensation, highlighting key research challenges. Section III introduces the proposed methodology. Section IV presents experimental validations and performance analysis, and Section V concludes the paper with a summary and future research directions.
@@ II State-of-the-art and research challenges
Shape deviation modeling and compensation are critical for improving geometric accuracy and quality control in AM. Early work focused on analytical and statistical models, while recent advances leverage data-driven methods such as machine learning. This section highlights three key areas: (1) Deviation Prediction Modeling, (2) Deviation Compensation Modeling, and (3) Research Challenges and Our Proposed Methodology, emphasizing the shift to data-driven strategies and the hurdles in achieving industrial-scale precision in AM mass production.
@@ II-A Deviation Prediction Modeling
Early research on shape deviation statistical and analytical parameterization models in 2D and 3D[ZHU2019496, xu2017reverse, Afazov02012021]. Schmutzler et al. [schmutzler2016compensating] framed deviation prediction as a non-rigid shape registration problem, using B-spline registration to predict deviations and compensate CAD parts by adjusting vertices oppositely [luan2017statistical]. Huang et al. [huang2014statistical] used Polar and Spherical Coordinate Systems for in-plane and out-of-plane deviation prediction, though their methods required extensive calibration, limiting scalability and adaptability. Finite element analysis (FEA) was also used to simulate deformation based on process parameters [10.1115/1.4041626].
More recently, machine learning (ML) techniques have been incorporated into deviation prediction to improve accuracy. Ferreira et al. [ferreira2019automated] used Bayesian neural networks to model and compensate for deviations. Decker et al. [decker2021geometric] developed a random forest model for 3D freeform shapes, enabling rapid, minimal human intervention. Convolutional neural networks (CNNs) [10015471] and Shape Deviation Generators [8957379] have been applied to detect complex deviation patterns, while Wang et al. [9786748] extended ML to both smooth and non-smooth shapes. Li et al. [LI2021101695] proposed in-situ deviation monitoring during printing.
Despite recent advances, several challenges persist. Analytical and statistical models remain computationally intensive, with simulations for a single part often requiring hours. Although machine learning approaches demonstrate potential, they are frequently restricted to simpler geometries—such as cylinders and polyhedrons—leaving a gap in effectively handling more complex 3D shapes. This underscores the need for more efficient, generalized models that can address both computational demands and the geometric complexity encountered in practical additive manufacturing applications.
@@ II-B Deviation Compensation Modeling
Input file modification methods have emerged as an alternative to directly correct deviations at the STL file level. Techniques like the Vertex Translation Algorithm (VTA) [navangul2013error] and the Surface-based Modification Algorithm (SMA) [zha2015geometric] iteratively refine geometric accuracy by minimizing chordal and staircase errors. However, these methods are computationally heavy, often increasing STL file sizes and requiring several iterations to meet tolerance standards. Afazov et al. [AFAZOV2017269] proposed a method to reduce residual stresses and distortions by interpolating 3D scan data, using a mathematical model to reverse distortions and pre-distort the CAD model. Their work demonstartes the effectiveness in reducing distortion on various parts printed with laser powder bed fusion from $\pm$300\text{\,}\mu\mathrm{m}$\$ to $\pm 65\text{\,}\mu\mathrm{m}$ micron tolerance. In a further development [AFAZOV201715], they integrated finite element analysis (FEA) with a thermal model for selective laser melting (SLM), although the computational runtime limited its industrial applicability.
To address these challenges, advanced algorithms have been developed. Bayesian models [ferreira2019automated] leverage uncertainty quantification to correct geometric errors in stereolithography printing, improving robustness across different printers. Decker et al. [decker2021geometric] employed random forest models to predict deviations in fused deposition modeling (FDM) using small training datasets, achieving over 44% shape deviation reduction, though requiring geometry-specific predictor variables. Hong et al. [HONG2021101594] applied artificial neural networks (ANN) to enhance dimensional accuracy in selective laser melting of truss lattice structures, demonstrating the potential of data-driven approaches for flexible freeform geometries and improved compensation. Despite these advances, most compensation methods focus on individual shapes, neglecting the significant thermal effects that vary with part position in the build chamber. Modeling these position-dependent deviations demands extensive resources, even for a single geometry.
@@ II-C Research Challenges and Proposed Methodology
Despite significant progress in shape deviation modeling and compensation, major challenges remain in scalability, data dependency, and generalization across varied geometries and process conditions. Addressing these gaps is critical to improving the efficiency and accuracy of additive manufacturing (AM). This work proposes a computational framework that predicts and compensates for geometric deviations by explicitly incorporating position-specific factors, such as spatial energy variations within the build chamber, for the same geometry. Given the limited prior research on position-dependent compensation in AM, this contribution is especially timely and impactful. The key challenges in this context include:
@@
  • Generalizability Across Complex Printing Geometries: Existing works on shape error compensation have demonstrated success in modeling, predicting, and compensating deviations for simple or structured geometries, such as 2D shapes or basic 3D forms like domes and stacked rectangles [8256230, 3106707f09f04e2982a488210eaf1bdb, 8957379, 9786748]. However, these methods suffer from several key limitations that prevent generalizability to arbitrary, non-smooth 3D freeform shapes. Finite Element Analysis (FEA)-based models are computationally expensive and impractical for complex industrial geometries and batch processes [jmmp4040112]. More recent approaches, utilizing neural networks [ZHU2020534, inproceedings_shen], preprocess geometries into fixed-size images or binary formats, which limits their resolution and results in a loss of detail.
  • Adaptability to Position-Dependent Batch Production: Traditional and machine learning-based compensation models often fail to consider the dynamic thermal and mechanical factors that influence AM, leading to suboptimal results under varying conditions [MCGREGOR2021101924]. Thermal-induced deformations significantly impact part distortion. Research by Chen et al. [10383123] shows how thermal variations across the print bed result in significant differences in part quality. For example, Figure 2 illustrates a dataset consisting of 140 identical bars placed at different positions within the printing bed. The shape deviation patterns exhibit significant variation, with parts positioned at the edges of the print chamber, such as Part ID 140 and Part ID 31 in Fig. 2 (b), experiencing greater deviations compared to those located at the center, such as Part ID 72 and Part ID 24. Hartmann et al. [technologies7040083] highlighted the importance of batch production in industrial settings, demonstrating a data-driven compensation approach applied to stacked fin-shaped parts. However, their method is constrained by the need for iterative redesigns and was validated using only a single geometry.
@@
      DGCNN (Dynamic Graph Convolutional Neural Network): To further enhance the model’s ability to handle the complexities of point cloud data, we incorporate DGCNN[wang2019dynamic] as the backbone architecture. DGCNN dynamically constructs graph structures by connecting points based on local geometric relationships, allowing it to effectively capture the varying densities and topologies inherent in complex shapes. The Edge Convolution (EdgeConv) operation—defined as a function that aggregates features from both a point and its neighbors by combining their relative and absolute positions—adapts to these local variations, facilitating effective feature extraction across the graph. This enables the model to achieve higher performance in tasks that require precise understanding of local geometry and complex spatial relationships, crucial for accurate shape prediction and deviation analysis in diverse 3D structures.
@@
      Incorporation of Part Position Information: Unlike standard shape classification or segmentation tasks—where invariance to global position is often desirable—our problem formulation relies heavily on the absolute spatial location of each part within the build volume. Thermal gradients, heat dissipation patterns, and residual stress accumulation are all influenced by a part’s position on the build plate. To capture these effects, we explicitly encode part placement information in the input data. Specifically, for parts translated along the x and y axes, we adjust the corresponding point cloud coordinates to account for translational offsets due to different positions on the build plate. For parts that are rotated, we apply a rotation matrix to the point cloud coordinates. This preprocessing step preserves both the intrinsic geometry of each part and its global location within the build chamber. By embedding this positional context into the model input, we enable the network to learn spatially dependent deformation behaviors, improving predictive accuracy under varying thermal and process conditions.
@@ III Methods
Review of Geometric Deep Learning. Convolutional Neural Networks (CNNs) have driven major advances in computer vision and NLP, but their application to 3D design data is limited due to the non-Euclidean nature of such structures. Graph Neural Networks (GNNs) offer a natural alternative by operating on graph-structured data $G={\mathcal{V},\mathcal{E}}$, where $\mathcal{V}$ and $\mathcal{E}$ represent vertices and edges, respectively. This framework allows for modeling of irregular geometries in CAD models and point clouds.
Graph neural networks can be broadly categorized into two families: spectral and spatial. Spectral methods define convolutions in the frequency domain using graph Laplacian eigen-decomposition [bruna2013spectral], though they are computationally intensive and less flexible for large-scale or dynamic graphs. Approximations using polynomial filters have been proposed to address this [levie2018cayleynets]. In contrast, spatial GNNs perform convolutions through localized message passing in the node domain, making them more scalable and suited for complex, irregular graphs [monti2017geometric]. A typical layer updates each node’s features by aggregating information from its neighbors:
@@
Here, $x_{i}^{(l)}$ denotes the representation of node $i$ at convolutions layer $l$, $h_{\theta}$ is a learnable aggregation function. Each node updates its representation by aggregating features from its immediate neighbors, taking into account both node attributes and edge information $e_{ij}\in\mathcal{E}$. :
Prominent spatial GNN architectures include GraphSAGE[DBLP:journals/corr/HamiltonYL17], Graph Attention Networks (GAT)[veličković2018graphattentionnetworks], and various Graph Convolutional Network (GCN) variants. EdgeConv introduces an edge-aware operation by concatenating each node’s feature $x_{i}^{(l)}$ and $x_{i}^{l}-x_{j}^{(l)}$, representing edge information with relative distance between neighboring vertices [wang2019dynamic]. MoNet [monti2017geometric] projects neighboring points onto a pseudo-coordinate plane, then define convolution operations on projected points on the plane. Overall, the evolution of graph neural networks presents a promising path for tackling geometric applications, offering flexibility and scalability in handling non-Euclidean data structures inherent in 3D designs [Chen_2024].
In our proposed framework, we adopt a spatial GNN due to its effectiveness in capturing localized deformation patterns and flexibility in handling varying mesh topologies. While spectral methods were considered, their computational cost and limited adaptability to local spatial variation make them less practical for our application.
Figure 4: The procedural sequence of the DL prediction engine begins with the conversion of the input CAD into an isometric mesh ($\mathcal{C}_{i}=\{\mathbf{c}_{1},\dots,\mathbf{c}_{k}\}$). Subsequently, graph neural networks are applied to predict shape deviation ($\hat{\mathcal{S}}_{i}=\{\hat{\mathbf{s}}_{1},\dots,\hat{\mathbf{s}}_{k}\}$). The accuracy of the predicted shape deviation is assessed by comparing it with the printed and scanned shapes ($S_{i}=\{\mathbf{s}_{1},\dots,\mathbf{s}_{k}\}$), utilizing a deformation loss function defined in Equation (7).
@@
Here, $G$ and $D$ represent the generator and discriminator, respectively. While the optimization problem above is equivalent to minimizing the Jensen-Shannon divergence (JSD) [goodfellow2014generative] when the discriminator is perfectly trained, JSD often leads to unstable GAN training, especially in the presence of singular measures [arjovsky2017wasserstein]. To address this issue, Wasserstein-GAN was introduced, replacing JSD with the Wasserstein metric, which compares transportation costs between distributions $\mathbb{P}_{G}$ and $\mathbb{P}_{r}$. The 1-Wasserstein distance derived from Kantorovich-Rubenstein duality is employed for GAN optimization, instead of directly computing transportation mappings. To ensure the computation of 1-Wasserstein distance, both the generator and discriminator must exhibit 1-Lipschitz continuity during training. This can be achieved through techniques such as weight clipping [arjovsky2017wasserstein] or gradient penalty.